# A Unified Framework for Entropy Search and Expected Improvement in Bayesian Optimization

### v1

@@ 2.2 Acquisition Functions
Various acquisition functions (AFs) have been proposed to balance exploration and exploitation in optimization tasks, each tailored to different problem characteristics and assumptions. These include Probability of Improvement (PI), Expected Improvement (EI) (Mockus, 1998), Upper Confidence Bound (UCB) (Srinivas et al., 2010), Knowledge Gradient (KG) (Frazier et al., 2008), and information-theoretic AFs (Villemonteix et al., 2009; Hennig & Schuler, 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017; Hvarfner et al., 2022; Tu et al., 2022). Below, we discuss two types of acquisition functions relevant to this study.
@@ 2.3 Related Work
It is widely recognized that EI can be prone to over-exploitation (Qin et al., 2017; Berk et al., 2019; De Ath et al., 2021). To mitigate this issue, Hoffman et al. (2011) and Kandasamy et al. (2020) propose to use a portfolio of AFs, which assigns probabilities to different AFs at each step. Another approach is Weighted EI (WEI), which adaptively adjusts the weights of the components within the EI acquisition function (Sóbester et al., 2005; Benjamins et al., 2023). Similarly, Qin et al. (2017) suggest “weakening” EI using sub-optimal points suggested by the AF to mitigate its over-exploitative behavior. However, these methods are primarily based on heuristics. Furthermore, information-theoretic acquisition functions are often excluded from these design enhancements, as they are generally considered distinct from heuristic AFs such as PI, EI, UCB, or KG.
@@ 3.1 Entropy Search Lower Bound
The idea behind our Variational Entropy Search (VES) framework is to maximize a variational lower bound of MES with a predetermined family of densities to approximate $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$. Since we assume noiseless observations, the support is $[\max\{y_{\bm{x}},y_{t}^{*}\},+\infty)$. VES is illustrated in Fig. 1. The lower bound is formalized in Theorem 3.1 and proven in Appendix A.1.
@@ 3.2 EI Through the Lens of the VES Framework
Theorem 3.2 reveals that EI can be viewed as a special case of MES, giving a new information-theoretic interpretation of the most popular acquisition function in use today. However, the exponential distribution has a fairly rigid parametric form that does not capture the characteristics of $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$. Fig. 1 (right) shows an example of the structural limitations of the exponential density in green. We generate $1000$ samples from an example distribution $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$, and observe that it significantly deviates from an exponential distribution. Specifically, the density of $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$ is non-monotonic, exhibiting a peak before decreasing near $\max\{y_{\bm{x}},y^{*}_{t}\}$ (approximately 1.55), while exponential distributions are necessarily monotonic.
This observation motivates the need to enrich the variational distributions $\mathcal{Q}$ to allow more flexibility. A natural extension is to use a Gamma distribution, which is a generalization of the exponential distribution. The Gamma density approximation in the previous example is shown in red in Fig. 1 (right). The next section introduces VES-Gamma, which is a more general AF that extends VES-Exp and its equivalent EI acquisition function.
@@ 3.3 VES-Gamma: A Generalization of EI
The ESLBO in Eq. (12) serves as the primary objective in the VES-Gamma algorithm. Eq. (12) consists of five terms, with the last term being the EI AF in Eq. (1) scaled by a multiplicative factor. The two hyperparameters, $k$ and $\beta$, originally part of the Gamma distribution, dynamically balance different components of the objective. In particular, when $k=1$, the Gamma distribution reduces to an exponential distribution, making the ESLBO in Eq. (12) equivalent to Eq. (10). In the following section, we discuss the approach for determining optimal values for $k$ and $\beta$.
@@
where $\psi(k)\coloneqq\partial\log\Gamma(k)/\partial k$ is the digamma function (Abramowitz et al., 1988), which can be efficiently approximated as a series. By Jensen’s inequality, $\log\mathbb{E}[z^{\ast}_{\bm{x}}]-\mathbb{E}[\log z^{\ast}_{\bm{x}}]\geq 0$. Since $\log k-\psi(k)$ is strictly decreasing and approaches zero asymptotically (see Fig. 2), the root of Eq. (13), $k^{\ast}_{\bm{x}}$, exists uniquely—except in the degenerate case where $z^{\ast}_{\bm{x}}$ is deterministic, in which case we apply a clamping function to prevent it from being zero.
@@
With this analysis, the optimal value $k_{\bm{x}}^{\ast}$ is determined by solving the root of Eq. (13) using a one-dimensional root-finding method, where expectations of $z^{\ast}_{\bm{x}}$ are estimated via Monte Carlo sampling from $p(y^{*},y_{\bm{x}}\mid\mathcal{D}_{t})$. Once $k^{\ast}_{\bm{x}}$ is obtained, the optimal $\beta^{\ast}_{\bm{x}}$ follows as:
@@
Implementing VES-Gamma in Algorithm 2 is computationally intensive. The number of inner iterations, $N$, must be sufficiently large for convergence, and each inner iteration requires estimating $\mathbb{E}[z^{\ast}_{\bm{x}}]$ by sampling a large number of $y^{*}$. Consequently, the overall BO loop takes significantly more time than EI and MES, as shown in Table 2 for $N=50$. However, since black-box function evaluations are often expensive, the additional computational cost of VES-Gamma is not a major bottleneck in many real-world applications.
Table 2: Average duration of a BO loop for each AF. We measure the runtime on the Branin, Levy, and Hartmann benchmarks and average over benchmarks, BO iterations, and 10 outer repetitions. VES has a considerably higher runtime than the other acquisition functions.
@@ 4.2 Comparing VES-Exp and EI
Several factors explain the remaining discrepancies between VES-Exp and EI. First, since both acquisition functions are non-convex, their optimization may yield different next sampling points $\bm{x}_{t+1}$ due to variations in initialization. Second, VES methods employ a clamping mechanism to ensure that $z^{\ast}_{\bm{x}}$ remains numerically positive, which introduces a dependency between $y^{\ast}$ and $\bm{x}$. In practice, this violates the assumptions used in the proof in Appendix A.2. Finally, while EI has a closed-form expression, VES-Exp relies on Monte Carlo estimation, introducing numerical inexactness and potential discrepancies.
Table 3: Kolmogorov-Smirnov two-sample test passing rate between VES-Exp and EI for various benchmarks.
@@ 4.3 Performance of VES-Gamma
Fig. 4 illustrates the performance of various methods, including MES, EI, and VES-Gamma, across four synthetic benchmark functions: Branin ($d=2$), Levy ($d=4$), Hartmann ($d=6$), and Griewank ($d=8$). The metric shown is the logarithm of the best value (or simple regret), averaged over 10 independent runs.
@@
Fig. 5 shows the optimization performance on the 100-dimensional GP prior samples. For $\ell=0.05,0.1,0.25$, VES-Gamma outperforms EI and MES by a wide margin. EI and MES converge to a suboptimal solution. Only for $\ell=0.5$ does EI reach the same quality as VES-Gamma, outperforming MES.
@@
Fig. 6 presents the performance of VES-Gamma, EI, and MES across four real-world optimization problems: the 60-dimensional Rover trajectory optimization, the 124-dimensional Mopta08 vehicle optimization, the 180-dimensional weighted Lasso-DNA regression, and the 388-dimensional SVM hyperparameter tuning benchmarks.
Consistent with previous observations, VES-Gamma delivers strong performance, significantly outperforming all other acquisition functions on the SVM benchmark. It also ranks among the top-performing methods, alongside EI, on the Mopta08 and Lasso-DNA benchmarks. On the Rover problem, VES-Gamma performs comparably to EI, while MES achieves the best results in this scenario.
@@ 5 Conclusion
Impact.
@@ A.2 VES-Exp and EI Algorithmic Equivalence
Theorem 3.2 is proved as follow:
@@ B.1 Synthetic Test Functions
Figure 7: Performance plots for EI, MES, VES-Gamma-Sequential, and VES-Gamma on additional synthetic benchmark functions. VES-Gamma shows robust performance throughout the bank.
Fig. 7 shows the performance of the different acquisition functions, EI, MES, VES-Gamma-Sequential, and VES-Gamma, on additional synthetic benchmark functions: the Ackley and Michalewicz test functions [^1] https://www.sfu.ca/~ssurjano/optimization.html and the Lasso-High and Lasso-Hard benchmarks (Šehić et al., 2022). On the 1000-dimensional Lasso-Hard problem, VES-Gamma ran into a timeout after 48 hours. Therefore, we plot the mean up to the minimum number of iterations performed across all repetitions. VES-Gamma demonstrates robust performance across the benchmarks, outperforming all other acquisition functions on Ackley, MES, and VES-Gamma-Sequential on Michalewicz, and performing similarly to the other acquisition functions on the Lasso benchmarks. VES-Gamma and MES perform considerably worse than VES-Gamma, especially on the more high-dimensional problems.

### v2

@@ 2.2 Acquisition Functions
Various acquisition functions (AFs) have been proposed to balance exploration and exploitation in optimization tasks, each tailored to different problem characteristics and assumptions. These include Probability of Improvement (PI), Expected Improvement (EI) (Mockus, 1998; Jones et al., 1998), Upper Confidence Bound (UCB) (Srinivas et al., 2010), Knowledge Gradient (KG) (Frazier et al., 2008), and information-theoretic AFs (Villemonteix et al., 2009; Hennig & Schuler, 2012; Hernández-Lobato et al., 2014; Wang & Jegelka, 2017; Hvarfner et al., 2022; Tu et al., 2022). Below, we discuss two types of acquisition functions relevant to this study.
@@ 2.3 Related Work
It is widely recognized that EI can be prone to over-exploitation (Qin et al., 2017; Berk et al., 2019; De Ath et al., 2021). To mitigate this issue, Hoffman et al. (2011) and Kandasamy et al. (2020) propose to use a portfolio of AFs, which assigns probabilities to different AFs at each step. Snoek et al. (2012) proposed a fully-Bayesian treatment on EI to improve empirical performance. Another approach is Weighted EI (WEI), which adaptively adjusts the weights of the components within the EI acquisition function (Sóbester et al., 2005; Benjamins et al., 2023). Similarly, Qin et al. (2017) suggest “weakening” EI using sub-optimal points suggested by the AF to mitigate its over-exploitative behavior. However, these methods are primarily based on heuristics. Furthermore, information-theoretic acquisition functions are often excluded from these design enhancements, as they are generally considered distinct from heuristic AFs such as PI, EI, UCB, or KG.
@@ 3.1 Entropy Search Lower Bound
The idea behind our Variational Entropy Search (VES) framework is to maximize a variational lower bound of MES with a predetermined family of densities to approximate $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$. Since we assume noiseless observations, the support is $[\max\{y_{\bm{x}},y_{t}^{*}\},+\infty)$. VES is illustrated in Figure 1. The lower bound is formalized in Theorem 3.1 and proven in Appendix A.1.
@@ 3.2 EI Through the Lens of the VES Framework
Theorem 3.2 reveals that EI can be viewed as a special case of MES, giving a new information-theoretic interpretation of the most popular acquisition function in use today. However, the exponential distribution has a fairly rigid parametric form that does not capture the characteristics of $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$. Figure 1 (right) shows an example of the structural limitations of the exponential density in green. We generate $1000$ samples from an example distribution $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$, and observe that it significantly deviates from an exponential distribution. Specifically, the density of $p(y^{*}\mid\mathcal{D}_{t},y_{\bm{x}})$ is non-monotonic, exhibiting a peak before decreasing near $\max\{y_{\bm{x}},y^{*}_{t}\}$ (approximately 1.55), while exponential distributions are necessarily monotonic.
This observation motivates the need to enrich the variational distributions $\mathcal{Q}$ to allow more flexibility. A natural extension is to use a Gamma distribution, which is a generalization of the exponential distribution. The Gamma density approximation in the previous example is shown in red in Figure 1 (right). The next section introduces VES-Gamma, which is a more general AF that extends VES-Exp and its equivalent EI acquisition function.
@@ 3.3 VES-Gamma: A Generalization of EI
The ESLBO in Eq. (12) serves as the primary objective in the VES-Gamma algorithm. Eq. (12) consists of five terms, with the last term being the EI AF in Eq. (1) scaled by a multiplicative factor. The two hyperparameters, $k$ and $\beta$, originally part of the Gamma distribution, dynamically balance different components of the objective. In particular, when $k=1$, the Gamma distribution reduces to an exponential distribution, making the ESLBO in Eq. (12) equivalent to Eq. (10). In the following section, we discuss the approach for determining values for $k$ and $\beta$.
@@
where $\psi(k)\coloneqq\partial\log\Gamma(k)/\partial k$ is the digamma function (Abramowitz et al., 1988), which can be efficiently approximated as a series. By Jensen’s inequality, $\log\mathbb{E}[z^{\ast}_{\bm{x}}]-\mathbb{E}[\log z^{\ast}_{\bm{x}}]\geq 0$. Since $\log k-\psi(k)$ is strictly decreasing and approaches zero asymptotically (see Figure 2), the root of Eq. (13), $k^{\ast}_{\bm{x}}$, exists uniquely—except in the degenerate case where $z^{\ast}_{\bm{x}}$ is deterministic. In the practical implementation, we apply a clamping function to ensure that the term $\log k-\psi(k)$ does not become zero, and we employ a regularization mechanism to keep the resulting root $k^{\ast}_{\bm{x}}$ close to 1. Specifically, we use L2 regularization when solving $\log k-\psi(k)=\log\mathbb{E}[z^{\ast}_{\bm{x}}]-\mathbb{E}[\log z^{\ast}_{\bm{x}}]$ since the unregularized version is unstable, presumably due to a widely flat landscape. In particular, for $\xi(k)\coloneqq\log k-\psi(k)-\log\mathbb{E}[z^{\ast}_{\bm{x}}]+\mathbb{E}[\log z^{\ast}_{\bm{x}}]$, we solve the following optimization problem:
$\min_{k}\xi(k)^{2}+\lambda\left(k-1\right)^{2},$ (14)
where $\lambda$ is a regularization parameter which is set to 1 in our experiments.
@@
With this analysis, the value $k_{\bm{x}}^{\ast}$ is determined by minimizing Eq. (14) using Brent’s method (Brent, 2013), where expectations of $z^{\ast}_{\bm{x}}$ are estimated via Monte Carlo sampling from $p(y^{*},y_{\bm{x}}\mid\mathcal{D}_{t})$. Once $k^{\ast}_{\bm{x}}$ is obtained, the corresponding $\beta^{\ast}_{\bm{x}}$ follows as:
@@
Implementing VES-Gamma in Algorithm 2 is computationally intensive. The number of inner iterations, $N$, must be sufficiently large for convergence, and each inner iteration requires estimating $\mathbb{E}[z^{\ast}_{\bm{x}}]$ by sampling a large number of $y^{*}$. Consequently, the overall BO loop takes significantly more time than EI and MES, as shown in Table 2 for $N=5$. However, since black-box function evaluations are often expensive, the additional computational cost of VES-Gamma is not a major bottleneck in many real-world applications.
Table 2: Average duration of a BO loop for each AF. We measure the runtime on the Branin, Levy, and Hartmann benchmarks and average over benchmarks, BO iterations, and 10 random restarts. For $N=5$ outer repetitions, VES has a higher runtime than the other acquisition functions.
@@ 4.2 Comparing VES-Exp and EI
Several factors explain the remaining discrepancies between VES-Exp and EI. First, since both acquisition functions are non-convex, their optimization may yield different next sampling points $\bm{x}_{t+1}$ due to variations in initialization. Second, VES methods employ a clamping mechanism to ensure that $z^{\ast}_{\bm{x}}$ remains numerically positive, which introduces a dependency between $y^{\ast}$ and $\bm{x}$. In practice, this violates the assumptions used in the proof in Appendix A.2. We also employed Log-EI (Ament et al., 2023) instead of EI in our experiment, which may also explain the difference. Finally, while EI has a closed-form expression, VES-Exp relies on Monte Carlo estimation, introducing numerical inexactness and potential discrepancies.
Table 3: Kolmogorov-Smirnov two-sample test passing rate between VES-Exp and EI for various benchmarks. More details about p-values are available in Figure 8 in the appendix.
@@ 4.3 Performance of VES-Gamma
Figure 4 illustrates the performance of various methods, including MES, EI, and VES-Gamma, across four synthetic benchmark functions: Branin ($d=2$), Levy ($d=4$), Hartmann ($d=6$), and Griewank ($d=8$). The metric shown is the logarithm of the best value (or simple regret), averaged over 10 independent runs.
@@
Figure 5 shows the optimization performance on the 100-dimensional GP prior samples. For $\ell=0.05,0.1,0.25$, VES-Gamma outperforms EI and MES by a wide margin. EI and MES converge to a suboptimal solution. Only for $\ell=0.5$ does EI reach the same quality as VES-Gamma, outperforming MES.
@@
Figure 6 presents the performance of VES-Gamma, EI, and MES across four real-world optimization problems: the 60-dimensional Rover trajectory optimization, the 124-dimensional Mopta08 vehicle optimization, the 180-dimensional weighted Lasso-DNA regression, and the 388-dimensional SVM hyperparameter tuning benchmarks.
Consistent with previous observations, VES-Gamma delivers strong performance, significantly outperforming all other acquisition functions on the SVM benchmark. It also ranks among the top-performing methods, alongside EI, on the Mopta08 and Lasso-DNA benchmarks. On the Rover problem, VES-Gamma performs comparably to EI, while MES achieves the best results in this scenario. MES exhibits mixed performance across the benchmarks, achieving the best results on Rover but falling behind on the Mopta08 and SVM problems.
@@ 5 Conclusion
Acknowledgements
This project was partly supported by the Wallenberg AI, Autonomous Systems, and Software program (WASP) funded by the Knut and Alice Wallenberg Foundation, the AFOSR awards FA9550-20-1-0138, with Dr. Fariba Fahroo as the program manager, DOE award DE-SC0023346, and by the US Department of Energy’s Wind Energy Technologies Office. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725
Impact Statement
@@ A.2 VES-Exp and EI Algorithmic Equivalence
Theorem 3.2 is proved as follows:
@@ B.1 Synthetic Test Functions
Figure 7: Performance plots for EI, MES, and VES-Gamma on additional synthetic benchmark functions. VES-Gamma shows robust performance throughout the bank.
Figure 7 shows the performance of the different acquisition functions, EI, MES, and VES-Gamma, on additional synthetic benchmark functions: the Ackley and Michalewicz test functions [^1] https://www.sfu.ca/~ssurjano/optimization.html and the Lasso-High and Lasso-Hard benchmarks (Šehić et al., 2022). On the 1000-dimensional Lasso-Hard problem, VES-Gamma ran into a timeout after 48 hours. Therefore, we plot the mean up to the minimum number of iterations performed across all repetitions. VES-Gamma demonstrates robust performance across the benchmarks, outperforming all other acquisition functions on Ackley, MES on Michalewicz, and performing similarly to the other acquisition functions on the Lasso benchmarks. VES-Gamma and MES perform considerably worse than VES-Gamma, especially on the more high-dimensional problems.
@@ Appendix C Kolmogorov-Smirnov Test Statistic
Detailed p-values for VES-Exp and EI Comparison.
We present the p-values obtained from the experiments detailed in Section 4.2. These results are illustrated in Figure 8. It is observed that for the majority of the sample pairs, the calculated p-values are substantially above the $5\%$ significance level.
Figure 8: Distribution of p-values for 500 sample pairs generated using the EI and VES-Exp acquisition functions.
Appendix D VES-Gamma Computational Acceleration
Table 2 highlights the higher computational cost of VES methods compared to EI and MES. However, we observe that a technique known as Variable Projection (VarPro) (Golub & Pereyra, 1973; Poon & Peyré, 2023) can be leveraged to accelerate the computation of VES under certain conditions, which VES-Gamma satisfies.
The key idea behind VarPro is that when the function ESLBO has a specific structure,
$\max_{\bm{x};k,\beta}\text{ESLBO}(\bm{x};k,\beta)=\max_{\bm{x}}\left(\underbrace{\max_{k,\beta}\text{ESLBO}(\bm{x};k,\beta)}_{\varphi(\bm{x})}\right),$ (18)
and the solution to $\max_{k,\beta}\text{ESLBO}(\bm{x};k,\beta)$ is unique, then $\varphi(\bm{x})$ is differentiable, with
$\frac{d}{d\bm{x}}\varphi(\bm{x})=\frac{\partial}{\partial\bm{x}}\text{ESLBO}(\bm{x},k^{\ast}_{\bm{x}},\beta^{\ast}_{\bm{x}}),$ (19)
where $k^{\ast}$ and $\beta^{\ast}$ are the unique values that maximize ESLBO.
Following the proof in Eq. (13), we establish that the solutions $k^{\ast}_{\bm{x}}$ and $\beta^{\ast}_{\bm{x}}$ are unique. This confirms that it is feasible to implement the VarPro strategy to accelerate the computation of VES-Gamma, eliminating the need for the iterative scheme in Algorithm 1. This ongoing work aims to reduce the computational cost of VES-Gamma to a level comparable to EI and MES.